OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x + y = 4 x - y = -6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (alexandervonhumboldt2):

x=4-y x=-6+y -6+y=4-y 2y=10 y=5 x=-1 -1, 5

OpenStudy (alexandervonhumboldt2):

C

OpenStudy (anonymous):

I have 4 more, can you help?

OpenStudy (anonymous):

@AlexandervonHumboldt2 4 More

OpenStudy (alexandervonhumboldt2):

ok but please post thm in separate post and tag me